Circular No. 6433

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995be IN ANONYMOUS GALAXY R. H. McNaught, Anglo-Australian Observatory, reports his discovery

of an apparent supernova on films taken on 1995 July 18 and Aug. 29 for the U.K. Schmidt Supernova Search by M. Hartley and M. J. Drinkwater; SN 1995be is present at the same brightness (mag ∼ 19) on both R films, but it is not present on films taken on 1996 Mar. 21, May 9, and 25, or on the second Palomar Sky Survey R photograph taken on 1992 Mar. 14. The position of SN 1995be from the Aug. 29 film is α = 15^{h}58^{m}14^{s}.30, δ = +19^{o}46′06′′.5 (equinox 2000.0), which is 16′′ east and 6′′ north of the galaxy’s center.

COMET C/1996 B2 (HYAKUTAKE) C. Lisse, M. Mumma, and R. Petre, NASA Goddard Space Flight

Center; and K. Dennerl, J. Englhauser, J. Schmitt, and J. Truemper, Max-Planck-Institut für Extraterrestrische Physik, Garching, communicate: “Comet C/1996 B2 was observed after its perihelion passage with the ROSAT High Resolution Imager from June 22.10 to 23.77 UT (when Δ = 1.16 AU, r = 1.35 AU, and visual m_{1} = 7.1) for 8900 s. The image, corrected for the proper motion of the comet, shows an extended source with a radial extent of at least 8′ (400 000 km). The peak surface brightness was at 0.002 count s^{−1} arcmin^{−2}, approximately a factor of 5 lower than during the previous ROSAT observations in March (IAUC 6373). The reacquisition of C/1996 B2 after perihelion confirms the continuous behavior of x-ray emission in the comet. ROSAT will continue to observe the comet until Sept. 8. The scheduled observation times will be posted at

URL http://www.rosat.mpe-garching.mpg.de/~jer/comets/; simulta-

neous observations at other wavelengths (especially extreme ultraviolet and radio) are encouraged.”

SS LACERTAE R. P. Stefanik, G. Torres, J. R. Caruso, and R. J. Davis, Harvard-

Smithsonian Center for Astrophyiscs, report: “An échelle spectrogram of SS Lac taken on July 6.20913 UT with the 1.5-m Wyeth Telescope at Oak Ridge Observatory clearly shows a double-lined spectrum, confirming the report on IAUC 6429. The velocities for the two components were +18 and −67 km/s.”

1996 July 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green